1) if you lean a ladder againist your house on level ground then how high will your ladder reach if it is 26 feet long and you set it 10 feet from the wall?
2)A cycle engine used a gas to oil mixture of 20:1. If you are to use 20 ounces of oil, then how many gallons of gas will you use (1 gallon is 128 ounces)?

Answer by jorel1380(2518)

(26)2-(10)2=n2
676-100=576=n2
n=24
The ladder will go 24 ft. high upon the wall.
-----------------------
20 oz. oil X 20:1=400 oz. of gas
400/128=3.125 gals of gas..